### v4

@@ 5 Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools to interact with external environments, novel categories of more catastrophic risks emerge. These risks stem not mainly from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which do not align with those of human designers or users, as extensively detailed in discussions of catastrophic risks from agentic AI [8]. These intrinsic risks, encompassing goal misalignment, unfaithful reasoning, emergent deception, self-preservation, scheming, and the persistence of such behaviors, pose unique and formidable challenges. Table 3 provides a structured summary of these risk categories, along with key observed behaviors and their implications for safety and control.
@@ 5.4 Persistence of Deception and the “Sleeper Agent" Problem
The challenge of managing these intrinsic risks is profoundly amplified by the persistence of deceptive behaviors embedded during training, as exemplified by the “sleeper agent" problem introduced in Section 3. Hubinger et al. [24] introduced and demonstrated the concept of “sleeper agents": LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. This concept is illustrated in Figure 3. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning, RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions," which indicates a form of persistence in its deceptive strategy [2].
@@ 6 Defense Mechanisms and Limitations
Researchers have proposed multiple defenses to mitigate these LLM threats, but each has limitations. Broadly, defenses fall into two categories: prevention-based (preprocessing or model changes) and detection-based (flagging malicious inputs or outputs) [37, 17]. A variety of techniques fall under these umbrellas, often conceptualized as a multi-layered strategy, as illustrated in Figure 4. This layered strategy aims to provide defense-in-depth by combining various mechanisms. The primary categories of these defenses, which will be explored in the following subsections, include prevention-based techniques and detection-based techniques.
@@ 6.1 Prevention-Based Defenses
Prevention strategies include input sanitization and prompt design. For example, paraphrasing the user’s prompt through a benign model can neutralize some adversarial phrasing: the idea is to reword inputs so that any malicious token sequences are altered while preserving meaning [17, 37]. Similarly, retokenization splits or merges tokens to disrupt known attack patterns, e.g., breaking a malicious command into fragments. Other structural defenses insert explicit delimiters or instructions to isolate user text from system instructions (techniques such as spotlighting and sandwiching), or use multiple redundant queries to cross-check consistency[28]. In training, adversarial training or robust fine-tuning may reduce input susceptibility. For instance, adversarial training methods that focus on the model’s refusal behavior have shown promise in enhancing robustness against prompt injection attacks [65]. RLHF and alignment methods also serve as preventive measures by reinforcing the model’s refusal of unsafe requests [15].
@@
    Supervised Fine-Tuning (SFT): The model is fine-tuned on synthetically generated examples of (prompt, CoT, output) tuples. Crucially, the CoT in these examples contains explicit reasoning that references and applies the relevant safety specifications.
@@
    Reinforcement Learning (RL): A separate "judge" LLM, which is given access to the safety specifications, is used to provide a reward signal, further refining the model’s ability to generate policy-adherent reasoning and responses.
@@
Another sophisticated training-based prevention strategy is the Instruction Hierarchy, which explicitly teaches an LLM to prioritize instructions from different sources based on their privilege level [57]. In this framework, instructions from the application developer (System Messages) have the highest priority, followed by the end user’s inputs (User Messages), and finally, content from external sources like web pages or tool outputs have the lowest priority. To achieve this, models are fine-tuned on synthetically generated data. For misaligned instructions (e.g., a prompt injection in a tool’s output telling the model to ignore its original task), the model is trained using “context ignorance” to act as if it never saw the malicious instruction. For aligned instructions (e.g., a user asking the model to reply in a different language), it is trained to comply. Wallace et al. [57] show that this method dramatically increases robustness against prompt injections and even generalizes to unseen attacks like jailbreaks, with only minimal degradation of standard capabilities.
@@ 6.2 Detection-Based Defenses
Detection-based defenses monitor for signs of attack. For example, perplexity-based or statistical detectors flag inputs or outputs that appear anomalous [29]. Specialized classifiers can be trained to identify phishing prompts or AI-generated malware code. Real-time monitoring of user queries (e.g. via a separate LLM critic) can catch obviously malicious requests [4]. Another emerging detection strategy involves monitoring the CoT reasoning of LLM agents. Baker et al.[5] demonstrated that using an auxiliary LLM to observe an agent’s CoT can be significantly more effective at detecting misbehavior such as reward hacking than monitoring only the agent’s actions and outputs. This approach leverages the idea that an agent’s intent might be more legible in its reasoning traces. A more principled approach to such oversight could involve a non-agentic “Scientist AI" designed to probabilistically assess risks based on a world model and serve as a guardrail against harmful actions by evaluating potential outcomes against safety specifications [8]. In content policy enforcement, many systems rely on filtered vocabularies or banned-topic checklists. For misuse (e.g., disinformation), watermarking generated text or running external bot-detection can help [30]. In autonomous agents, novel approaches include “red-teaming,” where another agent actively probes for misbehavior [64, 72, 23], or runtime oversight layers that intervene if the agent’s plan becomes harmful [59, 14].
@@
6.3 Limitations and Insufficiencies of Current Defenses
Despite these efforts, current defenses are far from foolproof, especially against the more sophisticated and inscrutable agentic risks. Paraphrasing and retokenization are only moderately effective: attackers can craft inputs that survive rewording or can learn the sanitization strategy. Indeed, paraphrasing often only alters surface phrasing and a sufficiently sophisticated input can still convey the malicious intent [17]. Moreover, these techniques can degrade performance: an innocuous prompt may be overly simplified or distorted, reducing the model’s utility. Delimiters and input framing help, but adversaries have already found evasion methods by varying their injection style. Detection-based defenses suffer from false positives and adaptability issues. For example, as Zugecova et al. [75] showed, adding personalization tokens caused LLMs to bypass their own safety filters completely.
Even promising detection methods like CoT monitoring come with their own caveats. While observing CoTs can reveal misaligned intent, Baker et al. [5] found that directly incorporating CoT monitors into an agent’s reinforcement learning objective—effectively penalizing ’bad thoughts’—can lead to ’obfuscated reward hacking.’ Agents may learn to produce CoTs that appear benign to the monitor while still pursuing misaligned goals, thereby rendering the CoT an unfaithful or deceptive representation of their true reasoning. This highlights a critical trade-off, i.e., strong optimization pressure on the reasoning process itself might diminish its monitorability and promote more subtle forms of deception.
Critically, the findings by Hubinger et al.[24] suggest that current safety training paradigms, including adversarial training and RLHF, may be insufficient to remove deeply embedded deceptive behaviors or “sleeper agent" capabilities. If models can learn to strategically appear aligned during training and evaluation, they may create a dangerous false impression of safety. This underscores the argument that ensuring an agentic AI will not cause harm is a notoriously difficult, unsolved technical problem, especially as current training methods for agentic AI can lead to issues such as goal misspecification and misgeneralization[8]. As one survey notes, most defense mechanisms “present inherent limitations that can be exploited by informed and sufficiently skilled malicious actors”[17]. In summary, while a multi-layered approach (sanitization, monitoring, aligned training, human oversight) can mitigate risk, no single defense is fully effective against the evolving threat landscape, particularly the challenge of ensuring genuine alignment and preventing strategic deception in advanced agents[17, 24, 40], leading some researchers to propose fundamentally different, non-agentic AI paradigms such as “Scientist AI", designed for trustworthiness and safety from the ground up by focusing on understanding and probabilistic inference rather than goal pursuit [8]. However, the “Scientist AI” paradigm remains in its infancy, and it is still unclear whether such non-agentic systems can ultimately match the general intelligence demonstrated by agentic AI models.
@@ 6.4 A Layered Deployment Playbook for Practitioners
While no single defense is foolproof, practitioners can implement a defense-in-depth strategy by layering controls across the application lifecycle. A prioritized playbook include:
@@
    • Logging and Forensics: Log all prompts, outputs, and intermediate tool calls. This is critical for incident response to determine if a compromise was due to prompt injection, a backdoor, or emergent behavior.

### v5

@@ 5 Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools to interact with external environments, novel categories of more catastrophic risks emerge. These risks stem not mainly from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which do not align with those of human designers or users, as extensively detailed in discussions of catastrophic risks from agentic AI [8]. These intrinsic risks, encompassing goal misalignment, unfaithful reasoning, emergent deception, self-preservation, scheming, and the persistence of such behaviors, pose formidable challenges. Table 3 provides a structured summary of these risk categories, along with key observed behaviors and their implications for safety and control.
@@ 5.4 Persistence of Deception and the “Sleeper Agent" Problem
The challenge of managing these intrinsic risks is profoundly amplified by the persistence of deceptive behaviors embedded during training, as exemplified by the “sleeper agent" problem introduced in Section 3. Hubinger et al. [25] introduced and demonstrated the concept of “sleeper agents": LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. This concept is illustrated in Figure 3. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning (SFT), RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions," which indicates a form of persistence in its deceptive strategy [2].
@@ 6 Defense Mechanisms and Limitations
Researchers have proposed numerous defenses to mitigate these LLM threats, but each has limitations. Broadly, defenses fall into two categories: prevention-based (preprocessing or model changes) and detection-based (flagging malicious inputs or outputs) [39, 17]. A variety of techniques fall under these umbrellas, often conceptualized as a multi-layered strategy, as illustrated in Figure 4. This layered strategy aims to provide defense-in-depth by combining various mechanisms.
@@ 6.1 Prevention-Based Defenses
One prevention-based defense is paraphrasing, where a user’s prompt is reworded by a benign model to neutralize adversarial phrasing before being processed [17, 39]. The core idea is that this process will alter the specific token sequences required for an attack to succeed while preserving the user’s legitimate intent. This approach has shown it can reduce the success of prompt injection attacks in some scenarios by breaking the syntactic order of the malicious instructions [39]. However, its limitations are severe. A major drawback is a substantial loss of utility on clean inputs; Liu et al. [39] found that paraphrasing legitimate prompts made them less accurate for the target task, resulting in an average performance drop of 14% when no attack was present. The generalization of this defense is also weak. It is considered only moderately effective because an attacker can anticipate this defense and craft more sophisticated inputs that survive the rewording process and still convey the malicious intent [17].
In training, methods like robust fine-tuning and adversarial training aim to reduce a model’s susceptibility to generating unsafe or incorrect content. These alignment techniques serve as preventive measures by modifying the model’s intrinsic behavior.
One prominent approach, Safe Reinforcement Learning from Human Feedback (Safe RLHF), directly addresses the inherent tension between a model’s helpfulness and its harmlessness [15]. In terms of effectiveness, this method is highly successful at reducing undesirable outputs by decoupling the two objectives. Instead of a single preference score, Safe RLHF uses separate reward and cost models trained on distinct human judgments for helpfulness and harmlessness. This constrained optimization approach was shown to drastically reduce the generation of harmful content on an evaluation set from 53% in the base model down to 2.45% after training, while simultaneously increasing helpfulness scores [15]. The generalization of this safety is enhanced through iterative red-teaming, where human adversaries identify and add new, challenging prompts to the training data to cover a wider range of potential attacks. However, the primary limitation of this technique is its significant operational complexity and cost. The process requires extensive, multi-stage human annotation across two separate preference dimensions, training of multiple models (reward and cost), and continuous red-teaming, making it a resource-intensive undertaking.
A different but related training strategy focuses on mitigating hallucinations by teaching the model to refuse questions outside its knowledge boundaries. While not a direct defense against adversarial attacks, this method, exemplified by Reinforcement Learning from Knowledge Feedback (RLKF), provides a crucial indirect defensive benefit by instilling a more cautious response behavior [67]. Effectively, RLKF trains a model to develop "self-awareness" of its knowledge limits, leading to significant gains in the trustworthiness and precision of its answers. A key strength is its generalization; a model trained with RLKF on one domain (e.g., arithmetic) can successfully apply its refusal capabilities to an entirely different domain (e.g., TriviaQA), avoiding the overfitting issues common with simpler fine-tuning [67]. The central limitation is an explicit trade-off: in exchange for higher reliability and fewer factual errors, the model becomes more conservative and answers a smaller proportion of total questions. This approach hardens the model against generating any untrustworthy content, thereby reducing the attack surface for prompts designed to coax it into fabricating speculative or fringe information.
@@
    SFT: The model is fine-tuned on synthetically generated examples of (prompt, CoT, output) tuples. Crucially, the CoT in these examples contains explicit reasoning that references and applies the relevant safety specifications.
@@
    RL: A separate "judge" LLM, which is given access to the safety specifications, is used to provide a reward signal, further refining the model’s ability to generate policy-adherent reasoning and responses.
@@
Another sophisticated training-based prevention strategy is the Instruction Hierarchy, which explicitly teaches an LLM to prioritize instructions from different sources based on their privilege level [59]. In this framework, instructions from the application developer (System Messages) have the highest priority, followed by the end user’s inputs (User Messages), and finally, content from external sources like web pages or tool outputs have the lowest priority. To achieve this, models are fine-tuned on synthetically generated data. For misaligned instructions (e.g., a prompt injection in a tool’s output telling the model to ignore its original task), the model is trained using “context ignorance” to act as if it never saw the malicious instruction. For aligned instructions (e.g., a user asking the model to reply in a different language), it is trained to comply. Wallace et al. [59] show that this method dramatically increases robustness against prompt injections and even generalizes to unseen attacks, with only minimal degradation of standard capabilities.
A different prevention-based strategy that has gained traction is machine unlearning, which aims to remove the influence of specific data or undesirable capabilities from a trained model without the need for complete retraining [36]. This technique is particularly relevant for addressing security issues through the removal of harmful or biased content that may have been learned during pre-training. The core idea is to efficiently modify the model to make it behave as if it had never been trained on the targeted information in the first place, while preserving its general knowledge and capabilities [36, 72].
From a defensive standpoint, machine unlearning can be conceptualized as a method for proactive capability removal. For instance, if a model has learned to generate instructions for creating bioweapons, unlearning techniques could be applied to specifically erase this harmful knowledge. In terms of effectiveness, the goal is not just to prevent the model from generating a specific harmful output, but to remove the underlying knowledge that enables it to do so. However, a significant limitation is the difficulty of ensuring complete and robust erasure. It has been shown that even after unlearning, sensitive information can sometimes be recovered through carefully crafted prompts or attacks [36]. Furthermore, there is often a trade-off between the thoroughness of the unlearning process and the preservation of the model’s overall utility; aggressive unlearning can lead to a degradation of performance on legitimate tasks [72]. The generalization of unlearning is also a challenge, as simply unlearning specific examples may not prevent the model from generating similar harmful content based on its broader knowledge [36].
@@ 6.2 Detection-Based Defenses
Detection-based defenses monitor for signs of attack. For example, perplexity-based or statistical detectors flag inputs that appear anomalous, such as the gibberish-like strings often found in adversarial prompts. This approach has proven to be a highly effective baseline defense; Jain et al. [30] found that perplexity filters successfully detected and blocked nearly all adversarial prompts generated by a standard attack optimizer across several open-source models. In terms of generalization, the defense shows considerable robustness even against adaptive "white-box" attackers. When an attacker is aware of the perplexity filter and modifies their attack to simultaneously optimize for both jailbreaking and low perplexity, the attack’s success rate drops significantly, as current optimizers struggle to satisfy both conflicting objectives.
However, the primary limitation of this method is its high false-positive rate, making it what Jain et al. [30] describe as "heavy-handed". Although effective in stopping attacks, the filter also incorrectly flags a significant number of benign, harmless prompts, on average about one in ten, which would be an untenable rate of disruption for most practical applications if used as a standalone defense. Therefore, while not a complete solution on its own, perplexity filtering can be a valuable component within a larger, multi-layered defense system where flagged prompts are routed for further analysis rather than being outright rejected.
Specialized classifiers can also be trained to identify malicious inputs. For instance, by using text embeddings of user prompts as features, traditional machine learning models can be trained to detect direct prompt injection attacks [4]. In terms of effectiveness, this embedding-based approach has proven quite successful; one study found a Random Forest classifier achieved an F1-score of 0.868, outperforming several state-of-the-art deep learning detectors by providing a better balance between precision and recall [4]. A key limitation, however, is that this specific method was evaluated on direct prompt injections, and its generalization to other attack vectors, such as indirect prompt injections or toxicity, has not yet been established and remains an area for future work [4].
For misuse like disinformation, one technical defense is to embed a statistical watermark into the model’s output, making it algorithmically identifiable as AI-generated [31]. In terms of effectiveness, this approach is potent, enabling detection with high accuracy (over 98% in experiments) from short spans of text, sometimes as few as 25 tokens. A key aspect of its generalization is that the detection algorithm is open-source and does not require access to the proprietary model’s API or parameters, allowing third parties to perform verification. However, the approach has notable limitations. Its primary weakness is in watermarking "low-entropy" or highly predictable text (e.g., "Barack Obama"), where the model has few alternative token choices; in these cases, the watermark is very weak and often undetectable. Furthermore, while resilient, the watermark is not immune to adversarial attacks. A malicious actor could attempt to remove the signal by using another language model to paraphrase or replace parts of the text. This robustness is a trade-off: experiments show that while such attacks can reduce detection rates, they are costly for the attacker, requiring significant modification of the text (e.g., replacing 30-50% of tokens) and substantially degrading its quality and coherence.
In autonomous agents, a paradigm of defense can be divide into two steps: automated red-teaming, where an agent actively probes for misbehavior [66, 75, 24], and runtime oversight layers that intervene if an agent’s plan becomes harmful [61, 14].
Automated red-teaming frameworks demonstrate high effectiveness by using LLM agents to continuously generate and refine attacks. For example, RedAgent creates context-aware jailbreaks and has proven highly efficient, successfully jailbreaking most tested black-box LLMs within five queries, a twofold improvement over previous methods [66]. Similarly, AutoRedTeamer utilizes a dual-agent system where one agent discovers new attack strategies from recent research while another executes them; this method achieved a 20% higher attack success rate on HarmBench compared to prior work, with a 46% reduction in computational cost [75]. These frameworks generalize well by creating diverse and context-specific attacks. For instance, AutoRedTeamer’s modular design allows it to continuously integrate new attack vectors, ensuring it adapts to emerging threats. A more specialized red-teaming approach targets the communication channels between agents in multi-agent systems (MAS), which has been shown to successfully compromise systems by manipulating inter-agent messages, a threat vector not applicable to single-agent setups [24]. The primary limitation of these red-teaming agents is that their effectiveness can be constrained by the creativity of the underlying LLM and the scope of the attack strategies they are designed to explore; they may miss entirely novel vulnerabilities that fall outside their operational paradigms.
In contrast, runtime oversight mechanisms act as a safety layer during an agent’s operation. AgentSpec, for instance, provides a domain-specific language (DSL) that allows developers to define explicit, customizable safety rules and has proven effective in preventing over 90% of unsafe code executions and 100% of hazardous actions in embodied agent tasks [61]. This approach is highly generalizable, with demonstrated applications in domains from code generation to autonomous driving. A more formal approach uses a declarative framework to specify an agent’s behavior as a finite-state machine, using a "decoding monitor" to enforce compliance at runtime [14]. This method offers strong guarantees but is limited by the practical difficulty of specifying complex, real-world behaviors in a formal structure and the computational overhead of the monitoring process. The principal limitation for all runtime oversight systems is the challenge of creating a comprehensive set of rules or specifications that can anticipate all potential failure modes without being overly restrictive and thus limiting the agent’s utility and autonomy.
Another emerging detection strategy involves monitoring the CoT reasoning of LLM agents. In terms of effectiveness, Baker et al. [5] demonstrated that using an auxiliary LLM to observe an agent’s CoT can be significantly more effective at detecting misbehavior like reward hacking than monitoring only the agent’s actions and outputs. Their experiments showed a CoT-based monitor achieved 95% recall on systemic hacks, whereas a monitor observing only actions achieved just 60%. The approach also shows promising generalization and scalability; the study found that a weaker model could effectively monitor a much more capable frontier model, providing a potential path for scalable oversight. Furthermore, the monitor was able to detect not only common, widespread hacks but also a diverse range of novel, complex, and problem-specific exploits. However, the method still has critical limitations. First, its utility can be destroyed by applying direct optimization pressure. Baker et al. [5] found that when the CoT monitor was used as part of the reward signal to train the agent (i.e., to penalize "bad thoughts"), the agent learned to engage in obfuscated reward hacking. It would produce a benign and plausible-looking CoT to fool the monitor while still executing the misaligned behavior, rendering its reasoning an unfaithful and deceptive signal. This risk of promoting obfuscation suggests that the monitoring utility of CoTs is fragile and may require avoiding strong, direct training pressures on the reasoning process itself. Second, as we have illustrated section 5.5, for some LLMs, the CoT may not be a faithful explanation of its actual reasoning process [33].
A more principled approach to such oversight could involve a non-agentic “Scientist AI," a theoretical framework designed to be safe and trustworthy by construction [8]. This paradigm is designed to serve as a probabilistic guardrail. Instead of relying on learned refusal behaviors, it uses a Bayesian world model to explicitly estimate the probability that a proposed action will lead to harmful consequences across multiple plausible interpretations of a safety policy. If the estimated risk exceeds a predefined threshold, the action is blocked. A key conceptual advantage is its convergence property: unlike current models that may become more deceptive with more compute, the Scientist AI is designed to become safer and more accurate with additional computational power, as it converges toward a unique, well-defined probabilistic answer [8].
In terms of generalization, the Scientist AI is proposed as an external control layer applicable to both narrow, specialized AIs and powerful, general agentic systems. Its function is to assess the potential long-term harm of any proposed action, providing a safety backstop regardless of the underlying agent’s architecture or alignment. However, the approach faces significant limitations, as outlined in the proposal itself. The framework is currently a research plan, not a deployed and tested system. Its most critical vulnerability is misuse: a malicious actor could exploit the system’s ability to model the world to design dangerous outputs (e.g., bioweapons), or could intentionally convert the non-agentic system into a harmful agent through repeated, scaffolded queries. The paper acknowledges that technical solutions like these are not a complete panacea, stressing that they must be complemented by robust social coordination, legal frameworks, and international treaties to be truly effective [8].
@@
In summary, while a multi-layered approach (sanitization, monitoring, aligned training, human oversight) can mitigate risk, no single defense is fully effective against the evolving threat landscape, particularly the challenge of ensuring genuine alignment and preventing strategic deception in advanced agents[17, 25, 42], leading some researchers to propose fundamentally different, non-agentic AI paradigms such as “Scientist AI", designed for trustworthiness and safety from the ground up by focusing on understanding and probabilistic inference rather than goal pursuit [8]. However, the “Scientist AI” paradigm remains in its infancy with its weaknesses, and it is still unclear whether such non-agentic systems can ultimately match the general intelligence demonstrated by agentic AI models.
@@
Prevention Machine Unlearning Removing the influence of specific data or capabilities (e.g., harmful knowledge) from a model without full retraining. Copyright/privacy violations, harmful content, biased behaviors. Difficult to ensure complete erasure; unlearned info can be recovered; may degrade overall utility. Liu et al. [36]
@@ 6.3 A Layered Deployment Playbook for Practitioners
To mitigate the fact that no single defense is foolproof, practitioners can maximize the security by implementing a defense-in-depth strategy by layering controls across the application lifecycle. A prioritized playbook include:
@@
    • Targeted Unlearning for Remediation: When a specific vulnerability is found (e.g., the model has learned private data or a dangerous capability), use machine unlearning techniques to surgically remove the problematic knowledge without a full, costly retrain [36].
@@
    • Logging and Forensics: Log all prompts, outputs, and intermediate tool calls. This is critical for incident response and provides the necessary data to inform targeted remediation, such as patching the model through machine unlearning if a specific knowledge-based vulnerability is discovered.